OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6

OpenStudy (texaschic101):

first we need to find the slope. Keeping in mind that in the formula y = mx + b, the slope is in the m position. Knowing this, in the equation y = 3x - 6, we find that the slope is 3. A parallel slope will be the same. Now we can use the same formula, y = mx + b, and sub in our slope and points and solve for b, the y intercept. slope(m) = 3 (2,-4) x = 2 and y = -4 now lets sub -4 = 3(2) + b -4 = 6 + b -4 - 6 = b -10 = b Therefore, your parallel line is : y = 3x - 10 Any questions ?

OpenStudy (anonymous):

No by looking at all the steps and seeing why u did it that way made me understand what the process is.thanks so much for ur help

OpenStudy (texaschic101):

no problem :)